OpenStudy (anonymous):

The area of the front face of the analog clock shown below is 113.04 square inches. The length of the minute hand is 0.15 inches less than the radius of the front face. What is the length of the arc, rounded to the nearest hundredth of an inch, the minute hand makes when it moves from the number 7 to the number 11 on the clock?

OpenStudy (anonymous):

@FoolForMath

OpenStudy (anonymous):

I am effete, tomorrow?

OpenStudy (anonymous):

We need to use the area formula first i think and that is= a*pi*r^2.. 113.04=3.14r^2... 113.04/3.14... I got that the radius is 6...and the length of the hand is 5.85... Then i think we have to use the CIRCUMFERENCE formula.... c=pi*d This has to be done now.. please try and help.. please

OpenStudy (anonymous):

I can't read now sorry :( Let me ask @ExperimentX.

OpenStudy (anonymous):

oh okay..... :( hopefully he can...

OpenStudy (experimentx):

4/12 * 2 * pi * r

OpenStudy (anonymous):

thats the formula?

OpenStudy (experimentx):

yeah ...

OpenStudy (anonymous):

so 12.56 is the answer??... THANKS:)

OpenStudy (experimentx):

must be ..

OpenStudy (anonymous):

do you think you can help me with one more question??

OpenStudy (experimentx):

if it's quite short

OpenStudy (anonymous):

okay! :)